# Dynamic Classification: Leveraging Self-Supervised Classification to Enhance Prediction Performance

### v1

@@ Abstract
In this paper, we propose an innovative dynamic classification algorithm designed to achieve the objective of zero missed detections and minimal false positives. The algorithm partitions the data into $N$ equivalent training subsets and $N$ prediction subsets using a supervised model, followed by independent predictions from $N$ separate predictive models. This enables each predictive model to operate within a smaller data range, thereby improving overall accuracy. Additionally, the algorithm leverages data generated through supervised learning to further refine prediction results, filtering out predictions that do not meet accuracy requirements without the need to introduce additional models. Experimental results demonstrate that, when data partitioning errors are minimal, the dynamic classification algorithm achieves exceptional performance with zero missed detections and minimal false positives, significantly outperforming existing model ensembles. Even in cases where classification errors are larger, the algorithm remains comparable to state-of-the-art models.
@@
The key innovations of this study include self-supervised classification learning, the use of small-range subset predictions, and the direct rejection of substandard predictions. While the current algorithm still has room for improvement in terms of automatic parameter tuning and classification model efficiency, it has demonstrated outstanding performance across multiple datasets. Future research will focus on optimizing the classification component to further enhance the algorithm’s robustness and adaptability.
@@
Key words: Dynamic Classification Algorithm, Prediction Accuracy Improvement, Self-supervised learning, Data Distribution, Subset Prediction, Industrial forecasts, Result Filtering
1 Background and Introduction
@@
The rapid increase in global energy demand has not only exacerbated environmental pollution but also intensified climate change and deepened the energy crisis. As a result, the development of renewable energy sources has become a crucial direction for achieving global sustainable development [1]. Currently, renewable energy sources such as solar, wind, and biomass are gradually becoming key alternatives to fossil fuels. However, as of 2017, fossil energy still accounted for 73.5% of the global electricity supply, while renewable energy made up only 26.5% [3]. Promoting the research, development, and application of new energy technologies has been recognized as a core strategy to address future energy shortages and environmental challenges.
@@
Energy storage technology, especially battery technology, plays a critical role in new energy systems. It not only promotes the efficient use of renewable energy but also plays a vital role in transforming energy structures [1]. Lithium-ion batteries, liquid current batteries, sodium-sulfur batteries, and other energy storage technologies have been rapidly developed in recent years, showing significant improvements in energy density, safety, and cost. However, there remains a gap in realizing the goal of "high safety, low cost, long lifespan, and environmental friendliness" [2]. Therefore, future technology development in energy storage is expected to feature the coexistence of various technological approaches.
@@
In recent years, with the in-depth study of battery materials, new battery technologies have emerged. Research has shown that optimizing material composition and battery design can significantly enhance battery performance, such as increasing energy density and cycle life [1]. Furthermore, the introduction of the circular economy concept has provided new directions for the sustainable development of batteries, such as efficient material recycling and reuse [2]. Nonetheless, battery technology still faces many challenges in achieving the goals of a fully sustainable lifecycle and environmental protection. Future research should focus on discovering new materials and the synergistic application of multiple technological approaches [1, 2].
@@
The application of artificial intelligence (AI) in the field of energy storage shows great potential. Using machine learning and big data technologies, AI accelerates the discovery, performance prediction, and optimal design of new materials. For instance, the materials genome program developed using AI has significantly improved the efficiency of battery materials development [4]. Additionally, AI has demonstrated strong capabilities in battery condition monitoring, fault diagnosis, and performance prediction, providing robust support for advancing battery technology [4]. In the future, AI is expected to further drive innovation and widespread adoption of battery technologies, injecting new vitality into energy storage development [4].
@@
The rapid advancement of AI technology offers a new opportunity for battery technology innovation. In recent years, machine learning algorithms have been widely applied in battery performance prediction and state-of-health (SOH) estimation, primarily focusing on non-invasive methods to obtain microstate information inside batteries [6]. For example, using data-driven machine learning models, charging and discharging curve characteristics can be analyzed to predict the capacity degradation and remaining lifetime of a battery [5]. Moreover, AI has accelerated the development of novel energy storage materials through high-throughput computing and material genome programs, significantly improving research efficiency and material screening accuracy [4].
@@
Regarding AI’s application in lithium-ion batteries, current research mainly focuses on model-based and data-based approaches [5]. Model-based methods analyze the chemical reactions and degradation mechanisms inside the battery. Though theoretically reflective of battery operations, these models often suffer from large errors due to complex internal chemical reactions and environmental factors [6]. Data-based methods, such as convolutional neural networks (CNN), recurrent neural networks (RNN), and support vector machines (SVM), have attracted attention for their flexibility and accuracy in nonlinear problems [5]. For instance, CNN and RNN models optimized by genetic algorithms (GA) have significantly improved battery capacity prediction, achieving a root mean square error (RMSE) of only 0.1176%, representing a considerable advancement over previous models.
@@
Despite these advances, challenges remain. The predictive performance of data-driven approaches relies heavily on high-quality data, which is costly and limited in quantity[6]. In addition, due to the different internal structures and operating conditions of different batteries, the generalization ability of the models is weak, making it difficult to adapt to complex and changing real-world situations[1]. Meanwhile, traditional methods usually rely on trial and error in parameter setting, a time-consuming and labor-intensive process that limits the practical application of these models[5]. Finally, there is a need in industry for a reliable method of picking out substandard products, not just a need to make predictions about a particular batch. In other words, there is an industrial need for 0 misses and fewer false positives. The current practice is basically to overlay the prediction with a classification model to pick out the substandard products, an approach that is difficult to strike a balance between misses and false kills, and is also unreliable.
@@
To address these challenges, this paper proposes a dynamic classification algorithm (DCA). The algorithm automatically divides the data into subsets by self-supervised learning and enables each subset to predict a small range of data based on the distributional features of the training set. Finally, the information from previous self-supervised learning is skillfully used to filter the predictions, which in turn reliably achieves the conditions of 0 misses and low false positives, thus eliminating the need for additional models. This approach significantly improves performance when the classification error is small, and rivals current state-of-the-art models when the classification error is large. Dynamic classification algorithms are highly adaptable and are expected to be widely used in various fields and industries.
@@ 2 Related Work
Numerous studies have explored methods to improve model prediction accuracy, with a focus on optimization algorithms, machine learning, and hybrid approaches. This section reviews key advancements in the field, highlighting their methodologies, strengths, and limitations, and provides a comparative analysis with the proposed Dynamic Classification Algorithm (DCA).
@@
Azevedo et al. [7] conducted a systematic review of hybrid optimization and machine learning methods for clustering and classification. Their approach integrated various techniques to overcome the limitations of single methods. However, their framework required significant computational resources and faced challenges in adapting to rapidly evolving data environments. Kotary et al. [8] proposed a joint prediction and optimization learning framework that directly learns optimal solutions from observable features. Despite its theoretical innovation, the approach demands end-to-end training for optimization problems, which can be computationally impractical, especially for nonconvex or discrete scenarios.
@@
Khan et al. [9] combined neural networks with traditional filters (e.g., Kalman and $\alpha$-$\beta$ filters) to enhance the prediction accuracy of dynamic systems under noisy conditions. While this method demonstrated significant improvements in dealing with uncertainty, it relied on complex parameter tuning, limiting its applicability in resource-constrained environments. Ippolito et al. [10] combined supervised and unsupervised learning to improve stratigraphic classification accuracy, offering probability distributions rather than discrete classifications. However, their method required extensive data preprocessing and feature engineering, which increased computational costs.
@@
Son et al. [11] applied both supervised and unsupervised learning to predict social media user engagement in the airline industry. Although innovative in its use of sentiment analysis for classification, the model required intensive data cleaning and customization for different business contexts. Shah and Satyanarayana [12] introduced a predictive range corrector model to enhance reliability by accounting for deviations between observed and predicted outputs. While this approach improved prediction accuracy, its reliance on control sequence optimization made it less scalable for broader applications.
@@
Dalal et al. [13] developed a supply chain optimization model using Convolutional Neural Networks (CNNs) and Bidirectional Long Short-Term Memory Networks (BiLSTMs). Their method improved sustainability by leveraging CNNs for resource allocation and BiLSTMs for temporal demand forecasting. Similarly, Guillén et al. [14] introduced the LSB-MAFS technique, combining Least Squares Boosting (LSB) and Multivariate Adaptive Regression Spline (MARS) to enhance production frontier forecasting. While effective, the method primarily addressed specific regression tasks, limiting its adaptability.
@@
Currently, data prediction results are often filtered through additional classification models. However, these models typically struggle to establish a reliable balance between missed detections and false positives in practical applications. This approach not only increases computational overhead but also risks over-filtering or omitting crucial data, thereby compromising the overall reliability of the prediction system. In contrast, the proposed Dynamic Classification Algorithm (DCA) employs self-supervised learning to automatically partition data subsets and, by predicting within small classification errors, directly filters results based on the distributional characteristics of the training set—without the need for additional models. The dynamic classification method ensures that the range for each subset is accurately defined, which allows the algorithm to achieve zero missed detections while effectively minimizing false positives, thus improving overall prediction accuracy. Furthermore, experimental results show that when the classification errors are minimal in the self-supervised subset partitioning phase, DCA significantly outperforms existing methods in terms of prediction stability and accuracy, and it demonstrates strong adaptability across various datasets and application scenarios.
@@ 4 Algorithm Structure Overview
The Dynamic Classification Algorithm (DCA) is designed to enhance prediction accuracy by partitioning the data into smaller, well-defined subsets. The algorithm consists of four core components, each contributing significantly to the overall effectiveness:
@@
  • Dynamic Classification Process – The core innovation of DCA, where self-supervised learning iteratively refines data segmentation, improving classification boundaries.
  • Redundant Training and Prediction – A mechanism that enhances stability by incorporating data from neighboring subsets, improving prediction accuracy.
  • Excluding Predicted Results – A filtering strategy that removes predictions failing to meet accuracy requirements, ensuring retained results achieve a balanced zero missed detections and minimal false positives.
This section provides a detailed breakdown of the algorithm’s structure, classification principles, redundant training mechanism, and the filtering of predicted results. The overall structure is illustrated in Fig.1, which divides the framework into four main components.
@@
The **dynamic classification module (red section)** is at the core of the DCA. The dataset is split into $Train_{t}$ and $Train_{p}$. Through iterative optimization, the algorithm adjusts classification boundaries based on the numerical distribution of training data, ensuring that each subset operates within a smaller, controlled data range instead of relying on broad, generalized classification.
The **redundant training and prediction module (blue section)** further optimizes classification by incorporating data from adjacent classification intervals. This module enhances model robustness by supplementing predictions with contextual data, reducing boundary inconsistencies.
The **excluding predicted results module (yellow section)** utilizes supervised learning insights to filter out unreliable predictions. By leveraging pre-learned segmentation boundaries, this module ensures that only high-confidence predictions are retained, effectively achieving a well-balanced zero missed detections and minimal false positives.
These four components collectively optimize classification, improve prediction accuracy, and enhance model adaptability across various datasets.
@@
4.1 Principle of Dynamic Classification Processes
The **Dynamic Classification Process** is a fundamental component of the Dynamic Classification Algorithm (DCA). In traditional classification models, the learning process typically begins by extracting feature representations, followed by clustering or classification. However, this approach can lead to the coexistence of high and low values within the same classification, a phenomenon that diminishes the predictive model’s ability to generalize effectively. Specifically, when target values span a wide numerical range, this inconsistency complicates the model’s learning process, as the features may be similar while the predicted values diverge significantly.
To address this issue, the DCA leverages **self-supervised learning** to learn from the distribution of predicted target values. The algorithm dynamically partitions the dataset into smaller, more homogenous subsets, iteratively adjusting classification boundaries until an optimal segmentation is found. This ensures that each subset contains target values within a narrower range, thereby improving prediction accuracy. The **Dynamic Classification Process** consists of six key modules: **Division Module**, **Classification Model Module**, **Convergence Judgment Module**, **Penalty Module**, **Correction Module**, and **Evaluation Module**, as illustrated in Fig.2.
Figure 2: Structure of the Dynamic Classification Algorithm
During the initialization phase, the training set is divided into $Train_{t}$ and $Train_{p}$ using a **1:1 ratio**, ensuring a balanced training and validation setup.
Division Module: The first step involves manually defining the number of planned classifications, **N**. The initial classifier applies a **Gaussian kernel function** to measure the distribution of values in the target column of $Train_{t}$, and the data is then divided based on the degree of fluctuation in the sorted values. Larger fluctuations indicate finer divisions, ensuring more precise segmentation (see Fig.3). This results in **N intervals**, defined by **N-1 cutoff points**, forming the initial segmentation list. This list is subsequently used in the Classification Model Module.
However, in cases where a specific accuracy requirement must be met, manual intervention can yield better results. For example, if only predictions achieving 98% or higher accuracy are considered valid, the predefined division ranges should align with this requirement. Considering a confidence interval of 95%, the division module can be manually adjusted so that the preset range aligns with $2\%\times 95\%$. This fine-tuning process enhances the precision of the classification boundaries, which in turn optimizes the performance of the Excluding Module.
@@ 4.1 Principle of Dynamic Classification Processes
Classification Model Module: This module trains a set of classification models using the $Train_{t}$ and $Train_{p}$ data. A variety of models—such as **Decision Trees**, **Random Forests**, and **LGBM** are employed, utilizing multi-threading to optimize performance. Each model is trained independently, with its performance evaluated through a confusion matrix on $Train_{p}$. The best-performing model is selected for the next iteration based on its ability to minimize classification errors.
Convergence Judgment Module: The confusion matrix output from the Classification Model Module is converted into percentages, with off-diagonal values representing the classification loss. The total loss for each iteration is recorded in **LossList**, and the model with the lowest loss (**BestLoss**) is identified as the **BestModel**. If the loss values do not change over **15 iterations**, or if the last **10 iterations** show significant instability, a warning is issued to prevent convergence to a local minimum.
Evaluation Module: To evaluate the model’s overall performance, a score is computed using Equation 4.1, which balances the best-case loss with the regular performance over the last 10 iterations. This scoring method ensures that models with occasional excellent performance but inconsistent overall results are not selected.
$\text{Score}=0.5\times\text{BestLoss}+0.5\times\frac{\sum_{i=n-10}^{n}a_{i}}{10},\quad a_{i}\in\text{LossList}$
Once all iterations are complete, the $OptimalClassificationModel$ and $OptimalSegmentationList$ are finalized.
In the event of non-convergence, the **Penalty Module** applies penalties based on two core factors:
@@
  • Number of Misclassified Samples: The more misclassified samples in a given class, the stronger the penalty. This is determined by counting the number of samples that do not fit within the target classification interval.
The penalty is calculated based on the misclassification degree. Larger deviations from the correct classification result in larger penalties. The penalty matrix is updated based on the following formula:
$Y=\left(\frac{X}{N}\right)^{2}+1$
where $X$ represents the misclassification degree, and $N$ is the number of classification intervals. This formula ensures that misclassifications with large deviations incur greater penalties.
After calculating the initial penalty, a secondary penalty is applied based on the number of misclassified samples. If a higher number of samples is misclassified in a particular class, the penalty increases. This penalty is calculated using the following formula:
$Y=X^{2}$
@@
Once the penalties are applied, the misclassified samples are reassessed. If the classification error exceeds an acceptable threshold, the model returns to the Classification Model Module for further iteration and refinement. This iterative process ensures that the model continues to improve until the penalty stabilizes at an acceptable level.

### v2

@@ Abstract
In this study, we propose an innovative dynamic classification algorithm aimed at achieving zero missed detections and minimal false positives, critical in safety-critical domains (e.g., medical diagnostics) where undetected cases risk severe outcomes. The algorithm partitions data in a self-supervised learning-generated way, which allows the model to learn from the training set to understand the data distribution and thereby divides training set and test set into N different subareas. The training and test subsets in the same subarea will have nearly the same boundary. For each subarea, there will be the same type of model, such as linear or random forest model, to predict the results of that subareas. In addition, the algorithm uses subareas boundary to refine predictions results and filter out substandard results without requiring additional models. This approach allows each model to operate within a smaller data range and remove the inaccurate prediction results, thereby improving overall accuracy. Experimental results show that, with minimal data partitioning errors, the algorithm achieves exceptional performance with zero missed detections and minimal false positives, outperforming existing ensembles like XGBoost or LGBM model. Even with larger classification errors, it remains comparable to that of state-of-the-art models.
@@
Key innovations include self-supervised classification learning, small-range subset predictions, and optimizing the prediction results and eliminate the unqualified ones without the need for additional model support. Although the algorithm still has room for improvement in automatic parameter tuning and efficiency, it demonstrates outstanding performance across multiple datasets. Future work will focus on optimizing the classification components to enhance robustness and adaptability.
@@
Keywords: Dynamic Classification Algorithm, Prediction Accuracy Improvement, Self-supervised learning, Data Distribution, Subset Prediction, Industrial forecasts, Result Excluding
1 Introduction
@@
The rapid increase in global energy demand has not only exacerbated environmental pollution but also intensified climate change and deepened the energy crisis. Consequently, the development of renewable energy sources has become a crucial direction for achieving sustainable global development [1]. Currently, renewable energy sources such as solar, wind, and biomass are gradually becoming key alternatives to fossil fuels. However, as of 2017, fossil energy still accounted for 73.5% of the global electricity supply, whereas renewable energy accounted for only 26.5% [3]. Promoting the research, development, and application of new energy technologies has been recognized as a core strategy for addressing future energy shortages and environmental challenges.
@@
Energy storage technology, particularly battery technology, plays a critical role in new energy systems. It not only promotes the efficient use of renewable energy but also plays a vital role in transforming energy structures [1]. Lithium-ion batteries, liquid current batteries, sodium-sulfur batteries, and other energy storage technologies have been rapidly developed in recent years, showing significant improvements in energy density, safety, and cost. However, there remains a gap in realizing the goal: high safety, low cost, long lifespan, and environmental friendliness [2]. Therefore, future technological developments in energy storage are expected to facilitate the coexistence of various approaches.
@@
In recent years, with an in-depth study of battery materials, new battery technologies have emerged. Research has shown that optimizing the material composition and battery design can significantly enhance battery performance by increasing the energy density and cycle life [1]. Furthermore, the introduction of the circular economy concept has provided new directions for the sustainable development of batteries, including efficient material recycling and reuse [2]. Nonetheless, battery technology still faces many challenges in achieving the goals of a fully sustainable lifecycle and environmental protection. Future research should focus on discovering new materials and the synergistic application of multiple technological approaches [1, 2].
@@
Artificial intelligence (AI) has significant potential in the field of energy storage. AI has accelerated the discovery, performance prediction, and optimal design of new materials using machine learning and big data technologies. For instance, a material genome program developed using AI has significantly improved the efficiency of battery materials development [4]. In addition, AI has demonstrated strong capabilities in battery condition monitoring, fault diagnosis, and performance prediction, providing robust support for advancing battery technology [4]. In the future, AI is expected to drive innovation and widespread adoption of battery technologies by, injecting new vitality into energy storage development [4].
@@
The rapid advancement of AI technology has offered new opportunities for innovation in battery technology. In recent years, machine learning algorithms have been widely applied in battery performance prediction and state-of-health (SOH) estimation, primarily focusing on noninvasive methods to obtain microstate information inside batteries [6]. For example, using data-driven machine learning models, the charging and discharging curve characteristics can be analyzed to predict the capacity degradation and remaining battery lifetime [5]. Moreover, AI has accelerated the development of novel energy storage materials through high-throughput computing and material genome programs, thereby significantly improving research efficiency and material screening accuracy [4].
@@
Regarding application of AI in lithium-ion batteries, current research mainly focuses on model-based and data-based approaches [5]. Model-based methods are used to analyze the chemical reactions and degradation mechanisms of the batteries. Although theoretically reflective of battery operations, these models often suffer from large errors due to complex internal chemical reactions and environmental factors [6]. Data-based methods, such as convolutional neural networks (CNN), recurrent neural networks (RNN), and support vector machines (SVM), have attracted attention owing to their flexibility and accuracy in nonlinear problems [5]. For instance, CNN and RNN models optimized by genetic algorithms (GA) have significantly improved battery capacity prediction, achieving a root mean square error (RMSE) of only 0.1176%, representing a considerable advancement over the previous models.
@@
Despite these advances, several challenges remain. The predictive performance of data-driven approaches relies heavily on high-quality data, which are costly and limited in quantity[6]. In addition, owing to the different internal structures and operating conditions of different batteries, the generalization ability of these models is weak, making it difficult to adapt to complex and changing real-world situations[1]. Traditional methods usually rely on trial and error in parameter setting, which is a time-consuming and labor-intensive process that limits the practical application of these models[5]. Finally, there is a need in the industry for a reliable method of picking out substandard products, and not just a need to make predictions about a particular batch. In other words, there is an industrial need for zero misses and fewer false positives. The current practice is to overlay the prediction with a classification model to pick out the substandard products, an approach that makes it difficult to strike a balance between misses and false kills, and is also unreliable.
In safety-critical domains, the stringent requirement for zero missed detection is intrinsically linked to system reliability and societal risk mitigation.In safety-critical domains, the stringent requirement for zero missed detection is intrinsically linked to system reliability and societal risk mitigation. Recent advances in self-supervised learning have demonstrated their potential in addressing distributional biases in industrial data. For instance, Zhang et al. [7] proposed a debiased contrastive learning framework that effectively reduces sampling bias in time-series fault detection, achieving an F1-score of 97.2%. Meanwhile, Wang et al. [8] developed a cognitive learning mechanism to dynamically adjust decision boundaries in multi-objective scenarios, showing a 7.58% improvement in task-specific accuracy. Building on these advancements, our dynamic classification algorithm integrates self-supervised data partitioning with adaptive boundary refinement to simultaneously achieve zero missed detections and low false positives across complex industrial scenarios. In autonomous driving, Chen et al. [9] demonstrated that perception models failing to detect pedestrians (with a missed detection rate exceeding $10^{-6}$) violate the ISO 21448 functional safety standards, potentially leading to fatal accidents. This finding echoes the foundational argument by Amodei et al. Similarly, in industrial manufacturing, Ren et al. [10] achieved zero-defect detection via deep learning, proving that a single missed defect could result in the scrapping of entire semiconductor batches, incurring multimillion-dollar economic losses. Collectively, these studies underscore that zero missed detection is not an aspirational goal but a non-negotiable constraint for safety-critical system design. Current technologies urgently require solutions such as the dynamic classification algorithm proposed in this work, which leverages self-supervised data partitioning and sub-domain model collaboration to simultaneously achieve zero missed detections and low false positives across complex scenarios—addressing cross-domain demands from autonomous driving to precision manufacturing.
@@
To address these challenges, this paper proposes a dynamic classification algorithm (DCA). The algorithm automatically divides the data into subsets using self-supervised learning and enables each subset to predict a small range of data based on the distributional features of the training set. Finally, the information from previous self-supervised learning is skillfully used to filter the predictions, which in turn reliably achieves the conditions of 0 misses and low false positives, thus eliminating the need for additional models. This approach significantly improves the performance when the classification error is small and rivals the current state-of-the-art models when the classification error is large. Dynamic classification algorithms are highly adaptable and expected to be widely used in various fields and industries.
@@ 2 Related Work
Numerous studies have explored methods to improve the model prediction accuracy, with a focus on optimization algorithms, machine learning, and hybrid approaches. This section reviews the key advancements in the field, highlighting their methodologies, strengths, and limitations, and provides a comparative analysis with the proposed Dynamic Classification Algorithm (DCA).
@@
Azevedo et al. [11] conducted a systematic review of hybrid optimization and machine learning methods for clustering and classification. Their approach integrated various techniques to overcome the limitations of single methods. However, their framework requires significant computational resources and faces challenges when in adapting to rapidly evolving data environments. Kotary et al. [12] proposed a joint prediction and optimization learning framework that directly learns optimal solutions from observable features. Despite its theoretical innovation, this approach demands end-to-end training for optimization problems, which can be computationally impractical, especially for nonconvex or discrete scenarios.
@@
Khan et al. [13] combined neural networks with traditional filters (e.g., Kalman and $\alpha$-$\beta$ filters) to enhance the prediction accuracy of dynamic systems under noisy conditions. Although this method demonstrated significant improvements in dealing with uncertainty, it relied on complex parameter tuning, limiting its applicability in resource-constrained environments. Ippolito et al. [14] combined supervised and unsupervised learning to improve the stratigraphic classification accuracy, offering probability distributions rather than discrete classifications. However, their method requires extensive data preprocessing and feature engineering, which increases computational costs.
@@
Son et al. [15] applied both supervised and unsupervised learning to predict the social media user engagement in the airline industry. Although innovative in its use of sentiment analysis for classification, the model requires intensive data cleaning and customization in different business contexts. Shah and Satyanarayana [16] introduced a predictive range corrector model to enhance the reliability by accounting for deviations between the observed and predicted outputs. Although this approach improved prediction accuracy, its reliance on control sequence optimization made it less scalable for broader applications.
@@
Dalal et al. [17] developed a supply chain optimization model using Convolutional Neural Networks (CNNs) and Bidirectional Long Short-Term Memory Networks (BiLSTMs). Their method improves sustainability by leveraging CNNs for resource allocation and BiLSTMs for temporal demand forecasting. Similarly, Guillén et al. [18] introduced the LSB-MAFS technique, combining Least Squares Boosting (LSB) and Multivariate Adaptive Regression Spline (MARS) to enhance production frontier forecasting. Although effective, the method primarily addressed specific regression tasks, limiting its adaptability.
@@
Currently, the data prediction results are often filtered using additional classification models. However, in practical applications, these models typically struggle to establish a reliable balance between missed detections and false positives. This approach not only increases the computational overhead but also risks over-filtering or omitting crucial data, thereby compromising the overall reliability of the prediction system. In contrast, the proposed DCA employs self-supervised learning to automatically partition data subsets and, by predicting within small classification errors, directly filters results based on the distributional characteristics of the training set—without the need for additional models. The dynamic classification method ensures that the range for each subset is accurately defined, which allows the algorithm to achieve zero missed detections while effectively minimizing false positives, thereby improving the overall prediction accuracy. Furthermore, experimental results show that when the classification errors are minimal in the self-supervised subset partitioning phase, DCA significantly outperforms existing methods in terms of prediction stability and accuracy, and demonstrates strong adaptability across various datasets and application scenarios.
@@ 4 Algorithm Structure Overview
The DCA is designed to enhance prediction accuracy by partitioning the data into smaller, well-defined subsets. The algorithm consists of four core components, each of which contributes significantly to overall effectiveness:
@@
  • Dynamic Classification Process – The core innovation of DCA is that self-supervised learning iteratively refines data segmentation and improves classification boundaries.
  • Redundant Training and Prediction – A mechanism that enhances stability by incorporating data from neighboring subsets, improves prediction accuracy.
  • Excluding Predicted Results – A filtering strategy that removes predictions failing to meet accuracy requirements, ensuring that retained results achieve balanced zero missed detections and minimal false positives.
This section provides a detailed breakdown of the algorithm structure, classification principles, redundant training mechanism, and filtering of the predicted results. The overall structure is illustrated in Fig.1, which divides the framework into four main components.
Figure 1: Structure of Dynamic Classification Algorithm
@@
The **dynamic classification module (red section)** is at the core of DCA. The dataset is split into $Train_{t}$ and $Train_{p}$. Through iterative optimization, the algorithm adjusts classification boundaries based on the numerical distribution of training data, ensuring that each subset operates within a smaller, controlled data range instead of relying on broad, generalized classification. The boundary information of these subsets is used as pseudo-labels for subsequent processing.
The **redundant training and prediction module (blue section)** further optimizes classification by incorporating data from adjacent classification intervals. This module enhances model robustness by supplementing predictions with contextual data and reducing boundary inconsistencies.
The **excluding predicted results module (yellow section)** uses self-supervised learning pseudo-labels to filter out unreliable predictions. By leveraging pre-learned segmentation boundaries, this module ensures that only high-confidence predictions are retained, effectively achieving well-balanced zero missed detections and minimal false positives.
These four components collectively optimizing classification, improve prediction accuracy, and enhance model adaptability across various datasets. Given that the ”routine processing module” section is centered on fulfilling routine tasks and does not incorporate any novel content, it will be omitted during the subsequent introduction of the principle of the DCA.
4.1 Principle of Dynamic Classification Processes
The **Dynamic Classification Process** is a fundamental component of the DCA. In traditional classification models, the learning process typically begins with extracting feature representations, followed by clustering or classification. However, this approach can lead to the coexistence of high and low values within the same classification, which diminishes the predictive model’s ability to generalize effectively. Specifically, when target values span a wide numerical range, this inconsistency complicates thelearning process of the model, as the features may be similar while the predicted values diverge significantly.
To address this issue, DCA leverages **self-supervised learning** to learn from the distribution of the predicted target values. The algorithm dynamically partitions the dataset into smaller, more homogenous subsets and iteratively adjusts the classification boundaries until an optimal segmentation is found. This ensures that each subset contains target values within a narrower range, thereby improving prediction accuracy. The **Dynamic Classification Process** consists of six key modules: **Division Module**, **Classification Model Module**, **Convergence Judgment Module**, **Penalty Module**, **Correction Module**, and **Evaluation Module**, as illustrated in Fig.2.
@@ 4.1 Principle of Dynamic Classification Processes
During the initialization phase, the training set is divided into $Train_{t}$ and $Train_{p}$ using 1:1 ratio, ensuring a balanced training and validation setup.
Division Module: The first step involves manually defining the number of planned classifications, **N**. The initial classifier applies a **Gaussian kernel function** to measure the distribution of values in the target column of $Train_{t}$, and the data is then divided based on the degree of fluctuation in the sorted values. Larger fluctuations indicate finer divisions, ensuring more precise segmentation (see Fig.3). This results in **N intervals**, defined by **N-1 cutoff points**, forming the $InitialSegmentationList$. The **N-1 cutoff points** constitute the initial boundary, and the $InitialSegmentationList$ formed by them will be subsequently employed in the Classification Model Module. The final Boundary will be generated from the $OptimalSegmentationList$, which will be discussed later.
@@
However, in cases where a specific accuracy requirement must be met, manual intervention can yield better results. For example, if only predictions achieving 98% or higher accuracy are considered valid, the predefined division ranges should be aligned with this requirement. Considering a confidence interval of 95%, the division module can be manually adjusted such that the preset range aligns with 2% $\times$ 95%. This fine-tuning process enhances the precision of the classification boundaries, which in turn optimizes the performance of the Exclusion Module.
Classification Model Module: This module trains a set of classification models using the $Train_{t}$ and $Train_{p}$ data. A variety of models—such as **Decision Tree**, **Random Forest**, and **LGBM** are employed, utilizing multi-threading to optimize performance. Each model is trained independently, and its performance evaluated using a confusion matrix on $Train_{p}$. The best-performing model is selected for the next iteration based on its ability to minimize classification errors.
Convergence Judgment Module: The confusion matrix output from the Classification Model Module is converted into percentages, with off-diagonal values representing the classification loss. The total loss for each iteration is recorded as **LossList**, and the model with the lowest loss (BestLoss) is identified as the **BestModel**. If the loss values do not change over 15 iterations, or if the last 10 iterations show significant instability, a warning is issued to prevent convergence to a local minimum.
Evaluation Module: To evaluate the model’s overall performance, a score is computed using Equation 1, which balances the best-case loss with the regular performance over the last 10 iterations. This scoring method ensures that models with occasional excellent performance but inconsistent overall results are not selected.
$\text{Score}=0.5\times\text{BestLoss}+0.5\times\frac{\sum_{i=n-10}^{n}a_{i}}{10},\quad a_{i}\in\text{LossList}.$ (1)
The $OptimalClassificationModel$ and $OptimalSegmentationList$ are finalized once all iterations are complete.
In the event of non-convergence, the Penalty Module applies penalties based on two core factors:
@@
  • Number of Misclassified Samples: The greater the number of misclassified samples in a given class, the stronger the penalty. This is determined by counting the number of samples that do not fit the target classification interval.
Partition error stems from features lacking strong support or an unreasonable number of partition intervals. Sometimes, an overly large number of partitions can increase a considerable burden of classification. The penalty is calculated based on the degree of misclassification. Larger deviations from the correct classification resulted in larger penalties. The penalty matrix is updated based Equation (2):
$Y=\left(\frac{X}{N}\right)^{2}+1.$ (2)
where $X$ represents the degree of misclassification, and $N$ is the number of classification intervals. This formula ensures that misclassifications with large deviations incur higher penalties.
After calculating the initial penalty, a secondary penalty is applied based on the number of misclassified samples. If a larger number of samples are misclassified in a particular class, the penalty increases. This penalty is calculated using the Equation (2):
$Y=X^{2}$ (3)
@@
Once penalties are applied, the misclassified samples are reassessed. If the classification error exceeds an acceptable threshold, the model returns to the Classification Model Module for further iterations and refinement. This iterative process ensures that the model continues to improve until the penalty stabilizes at an acceptable level.